# Kramers nodal lines in intercalated TaS2 superconductors

### v1

@@ Results
The KNLs in noncentrosymmetric achiral little group have been clearly demonstrated in the exemplary case of the $P\bar{6}m2$ intercalated TMD superconductors consisting of InxTaS2 ($x$ = 1/2, 2/3, 1) and PbTaSe2, where In1/2TaS2 showcases the cleanest FS with a single KNL crossing the Fermi level, termed as the “ideal Kramers nodal line metal”. The spin-orbital texture in In1/2TaS2 directly observed via spin-resolved ARPES offers a natural explanation for the spin-valley polarization in the language of KNL, with a large SOC spin splitting up to around 250 meV. Furthermore in InTaS2, combined ARPES and angle-dependent QO studies point to the existence of the pinch points enforced by the KNLs, especially for the experimentally-probed closed $\epsilon$ pockets intersecting both the $\Gamma-M$ and $\Gamma-A$ KNLs. These pinch points are reminiscent of the 2D massless Dirac fermions on the surface of a 3D topological insulator, but reside on the intersection of the 3D FSs of an ideal KNL metal. Therefore, when the pinch points that harbor $n\pi$ ($n\in\mathbb{Z}$ and is an odd number) Berry phase are gapped by superconductivity, they produce nontrivial vortex spectra hosting chiral Majorana zero modes, which we elaborate in Supplementary Note 5 based on the model analysis of an ideal KNL. And in this work, both the KNLs with their associated pinch points and superconductivity are demonstrated in the InxTaS2 family. More importantly, our work establishes an entire large family of quantum materials as a platform for realizing and tuning KNLs – intercalated TMD compounds. Specifically, the In site can be populated with different concentrations of Tl, Pb, Bi, and Sn, the Ta site with Nb, and the S site with Se [32]. The intercalation not only produces the broken inversion symmetry – a central requirement for realizing KNLs, but can also be utilized to introduce correlated physics and electronic orders in the presence of KNLs, such as magnetic orders, charge-density waves, or superconductivity, a regime that has not been previously explored. Therefore, this material family offers promising ingredients for spin and valley transport, axionic quasiparticles, and topological superconductivity.
@@
Going beyond material selection, two other directions would deserve further exploration. The first one is the interplay between dimensionality and unconventional superconductivity. Recently, unconventional nodal superconductivity has been reported in monolayer 1H-TaS2 [52]. In the exfoliable intercalated $P\bar{6}m2$ TMD family represented by indium intercalated 1H-TaS2 single crystals, the nature of the superconductivity deserves further investigation, especially upon approaching the few-layer limit. The second direction is strong electronic correlations. As analyzed in Supplementary Note 5, due to the Dirac physics of the KNL model, the AHE can be driven by Zeeman or exchange fields along the KNLs. However, the strength of the anomalous Hall conductivity is inversely proportional to the velocity of the KNL band. In the case of InxTaS2, the dispersive KNLs would give rise to an AHE that is overwhelmed by the ordinary Hall contributions. One way to increase the anomalous Hall signal is to induce flat KNLs at the Fermi level, which can be achieved in localized $f$-orbital electronic systems with one example in $P\bar{6}m2$ being UAuSi [57]. Such scenario may open an alternative avenue for topological phases and observables driven by strong correlations.
@@ Acknowledgments
We thank Shiming Lei for discussion at the initial stage of the project. This work was mainly supported by the Department of Defense, Air Force Office of Scientific Research under Grant No. FA9550-21-1-0343. This research used resources of the Advanced Light Source and the Stanford Synchrotron Radiation Lightsource, both U.S. DOE Office of Science User Facilities under Contracts No. DE-AC02-05CH11231 and No. AC02-76SF00515, respectively. Part of the research also used resources of the National Synchrotron Light Source II, a U.S. Department of Energy (DOE) Office of Science User Facility operated for the DOE Office of Science by Brookhaven National Laboratory under Contract No. DE-SC0012704. The ARPES work at Rice University was also supported by the Gordon and Betty Moore Foundation’s EPiQS Initiative through grant No. GBMF9470. Y. Gao’s work at the National High Magnetic Field Laboratory was funded by the Gordon and Betty Moore Foundation’s EPIQS Initiative through ICAM-I2CAM, Grant GBMF5305. L. B. is supported by the US-DoE, BES program, through award DE-SC0002613. A portion of this work was performed at the National High Magnetic Field Laboratory, which is supported by National Science Foundation Cooperative Agreement No. DMR-2128556 and the State of Florida. Work at the University of California, Berkeley and Lawrence Berkeley National Laboratory was funded by the U.S. DOE, Office of Science, Office of Basic Science, Materials Sciences and Engineering Division under Contract No. DE-AC02-05CH11231 (Quantum Materials Program KC2202). We acknowledge support from the US National Science Foundation (NSF) Grant Number 2201516 under the Accelnet program of Office of International Science and Engineering (OISE). The KKR calculation work was supported by the project Quantum materials for applications in sustainable technologies (QM4ST), funded as project No. CZ.02.01.01/00/22_008/0004572 by P JAK, call Excellent Research. Research at UBC was undertaken thanks in part to funding from the Canada First Research Excellence Fund, Quantum Materials and Future Technologies Program. P.R.H. acknowledges funding from NSF grant no. DMR 2047193.
@@ Methods
$P=\frac{1}{S}\frac{I_{\uparrow}-I_{\downarrow}}{I_{\uparrow}+I_{\downarrow}},$
@@
$\delta P=P\cdot\sqrt{\frac{(\sqrt{I_{\uparrow}})^{2}+(\sqrt{I_{\downarrow}})^{2}}{(I_{\uparrow}+I_{\downarrow})^{2}}+\frac{(\sqrt{I_{\uparrow}})^{2}+(\sqrt{I_{\downarrow}})^{2}}{(I_{\uparrow}-I_{\downarrow})^{2}}},$
where the uncertainty of the spin-resolved photoelectron counts takes the form of $\sqrt{I_{\uparrow}}$ and $\sqrt{I_{\downarrow}}$ assuming the Gaussian statistics of $I_{\uparrow}$ and $I_{\downarrow}$. In this calculation, the uncertainty from the Sherman function is not taken into consideration. The large error bar on one data point in Fig. 3(e) comes from the close values of spin up and spin down counts in the raw data. By definition, identical spin up and down counts give rise to a divergent uncertainty of the spin polarization. Additionally, the positions of $k_{3}$ and $k_{4}$ in Fig. 3(f) were actually taken on the opposite sides of the $\bar{K}-\bar{M}-\bar{K^{\prime}}$, closer to the $\bar{\Gamma}$ hexagonal pocket than the cut in Fig. 3(a). We instead showed the cut to be outside of the BZ just for illustration purpose of a complete FS image. The curvature of the momentum cut on the FS in Fig. 3(a) and (f) is omitted to avoid curved and distorted band dispersions in the 3D plot.
ARPES data on InTaS2 and PbTaSe2 were taken at the Stanford Synchrotron Radiation Lightsource, Beamline 5-2 and the Brookhaven National Lab, National Synchrotron Light Source II, Beamline 21-ID, respectively. Both are equipped with a DA30 electron analyzer with vertical slit and have linear horizontal, linear vertical, and circularly polarized photons available. Only the data measured by linear horizontal light were included in the main text and the supplemental information. Samples of both kinds were cleaved in-situ under a based pressure better than $3\times 10^{-11}$ Torr and temperatures below 30K. All ARPES measurements in the standard mode maintain an energy and angular resolution superior to 20 meV and 0.1, while spin-ARPES on In1/2TaS2 has the energy and angular resolution better than 50 meV and 1.
@@ 0.7 Ab-initio calculations
The Bloch spectral function and the one-step model ARPES calculations were carried out using the spin-polarized relativisitic Korringa-Kohn-Rostoker (SPR-KKR) package [58], under the full-potential fully-relativistic four component Dirac formalism. Exchange-correlation potential within the local spin density approximation by Vosko, Wilk, and Nusair [59] was used. To model the random vacancy on the indium site in InxTaS2 ($x=$ 1/2, 2/3), single-site coherent potential approximation was employed to obtain an auxiliary effective medium that reproduces the concentration averaged scattering properties [60, 58]. The KKR equations were solved with an angular momentum cutoff of $l_{max}$ = 4 to account for the occupied 4f orbitals of Ta and the needs for spectroscopy calculations. The ARPES calculations considered a semi-infinite surface model terminated by the In atoms under the experimental geometry and used the layer-KKR multiple scattering theory, together with coherent potential approximation [61, 62]. Therefore, such theory takes into account all factors such as light polarization, matrix element, final-state, surface, disorder, relativistic, and multiple scattering effects. Experimental crystal structures of In1/2TaS2 and In2/3TaS2 were utilized in the calculations [32, 33].
@@
The data that support the findings of this study are available from the corresponding authors upon reasonable request.
@@
M.Y. and E.M. conceived the project. J.H. suggested the material family. Y.Gao synthesized the crystals and performed magnetotransport and magnetization measurements with the help from M.O. ,A.H., A.M., and L.B. under the supervision of E.M.. Y.Z., Y.Guo, Z.Y., and J.S.O. performed the ARPES experiments under the supervision of M.Y. and R.J.B, with the help from A.F., S.K.M, M.H., D.L., A.R., and E.V.. A.P. and Y.Z. performed the KKR calculations under the supervision of J.M.. X.G. carried out the wave-function based first-principles calculations, with the help from X.J.G., under the supervision of K.T.L.. P.H. provided theory support. Y.Z., Y.Gao, and A.P. contributed equally to this work. All authors contributed to the manuscript preparation.

### v2

@@ Discussion
The KNLs in noncentrosymmetric achiral little group have been clearly demonstrated in the exemplary case of the $P\bar{6}m2$ intercalated TMD superconductors consisting of InxTaS2 ($x$ = 1/2, 1) and PbTaSe2 (Supplementary Note 6), where In1/2TaS2 showcases the cleanest FS with a single KNL crossing the Fermi level, termed as the “ideal Kramers nodal line metal”. The spin-orbital texture in In1/2TaS2 directly observed via spin-resolved ARPES offers a natural explanation for the spin-valley polarization engendered by the underlying broken inversion symmetry of the KNL little group, with a large SOC spin splitting up to around 250 meV. Furthermore in InTaS2, combined ARPES and angle-dependent QO studies point to the existence of the pinch points enforced by the KNLs, especially for the experimentally-probed closed $\epsilon$ pockets intersecting both the $\Gamma-M$ and $\Gamma-A$ KNLs. These pinch points are reminiscent of the 2D massless Dirac fermions on the surface of a 3D topological insulator, but reside on the intersection of the 3D FSs of an ideal KNL metal. Therefore, when the pinch points that harbor $n\pi$ ($n\in\mathbb{Z}$ and is an odd number) Berry phase are gapped by superconductivity, they produce nontrivial vortex spectra hosting chiral Majorana zero modes, which we elaborate in Supplementary Note 8 based on the model analysis of an ideal KNL. And in this work, both the KNLs with their associated pinch points and superconductivity are demonstrated in the InxTaS2 family. More importantly, our work establishes an entire large family of quantum materials as a platform for realizing and tuning KNLs – intercalated TMD compounds. Specifically, the In site can be populated with different concentrations of Tl, Pb, Bi, and Sn, the Ta site with Nb, and the S site with Se [32]. The intercalation not only produces the broken inversion symmetry – a central requirement for realizing KNLs, but can also be utilized to introduce correlated physics and electronic orders in the presence of KNLs, such as magnetic orders, charge-density waves, or superconductivity, a regime that has not been previously explored. Therefore, this material family offers promising ingredients for spin and valley transport, axionic quasiparticles, and topological superconductivity.
@@
Going beyond material selection, two other directions would deserve further exploration. The first one is the interplay between dimensionality and unconventional superconductivity. Recently, unconventional nodal superconductivity has been reported in monolayer 1H-TaS2 [51]. In the exfoliable intercalated $P\bar{6}m2$ TMD family represented by indium intercalated 1H-TaS2 single crystals, the nature of the superconductivity deserves further investigation, especially upon approaching the few-layer limit. The second direction is strong electronic correlations. As analyzed in Supplementary Note 8, due to the Dirac physics of the KNL model, the AHE can be driven by Zeeman or exchange fields along the KNLs. However, the strength of the anomalous Hall conductivity is inversely proportional to the velocity of the KNL band. In the case of InxTaS2, the dispersive KNLs would give rise to an AHE that is overwhelmed by the ordinary Hall contributions. One way to increase the anomalous Hall signal is to induce flat KNLs at the Fermi level, which can be achieved in localized $f$-orbital electronic systems with one example in $P\bar{6}m2$ being UAuSi [56]. Such scenario may open an alternative avenue for topological phases and observables driven by strong correlations.
@@ Acknowledgments
We thank Shiming Lei for discussion at the initial stage of the project. This work was mainly supported by the Department of Defense, Air Force Office of Scientific Research under Grant No. FA9550-21-1-0343. This research used resources of the Advanced Light Source and the Stanford Synchrotron Radiation Lightsource, both U.S. Department of Energy, Office of Science User Facilities under Contracts No. DEAC02-05CH11231 and No. DE-AC02-76SF00515, respectively. Part of the research also used resources of the National Synchrotron Light Source II, a U.S. Department of Energy (DOE) Office of Science User Facility operated for the DOE Office of Science by Brookhaven National Laboratory under Contract No. DE-SC0012704. The ARPES work at Rice University was also supported by the Gordon and Betty Moore Foundation’s EPiQS Initiative through grant No. GBMF9470 and the Robert A. Welch Foundation Grant No. C-2175 (M.Y.). Y. Gao’s work at the National High Magnetic Field Laboratory was funded by the Gordon and Betty Moore Foundation’s EPIQS Initiative through ICAM-I2CAM, Grant GBMF5305. L.B. is supported by the US-DoE, BES program, through award DE-SC0002613. A portion of this work was performed at the National High Magnetic Field Laboratory, which is supported by National Science Foundation Cooperative Agreement No. DMR-2128556 and the State of Florida. Work at the University of California, Berkeley and Lawrence Berkeley National Laboratory was funded by the U.S. DOE, Office of Science, Office of Basic Science, Materials Sciences and Engineering Division under Contract No. DE-AC02-05CH11231 (Quantum Materials Program KC2202). Y. Z. acknowledges support from the US National Science Foundation (NSF) Grant Number 2201516 under the Accelnet program of Office of International Science and Engineering (OISE). The KKR calculation work was supported by the project Quantum materials for applications in sustainable technologies (QM4ST), funded as project No. CZ.02.01.01/00/22_008/0004572 by P JAK, call Excellent Research. K. T. L. acknowledges the support of the Hong Kong Research Grant Council through Grants RFS2021-6S03, C6025-19G, C6053-23G, AoE/P- 701/20, 16307622, 16309223 and 16311424. Research at UBC was undertaken thanks in part to funding from the Canada First Research Excellence Fund, Quantum Materials and Future Technologies Program. P.H. acknowledges funding from NSF grant No. DMR 2047193. M.H. and D.L. acknowledge the support of the U.S. Department of Energy, Office of Science, Office of Basic Energy Sciences, Division of Material Sciences and Engineering, under Contract No. DE-AC02-76SF00515.
@@ Methods
$P=\frac{1}{S}\frac{I_{\uparrow}-I_{\downarrow}}{I_{\uparrow}+I_{\downarrow}},$ (1)
@@
$\delta P=P\cdot\sqrt{\frac{(\sqrt{I_{\uparrow}})^{2}+(\sqrt{I_{\downarrow}})^{2}}{(I_{\uparrow}+I_{\downarrow})^{2}}+\frac{(\sqrt{I_{\uparrow}})^{2}+(\sqrt{I_{\downarrow}})^{2}}{(I_{\uparrow}-I_{\downarrow})^{2}}},$ (2)
where the uncertainty of the spin-resolved photoelectron counts takes the form of $\sqrt{I_{\uparrow}}$ and $\sqrt{I_{\downarrow}}$ assuming the Poisson statistics of $I_{\uparrow}$ and $I_{\downarrow}$. In this calculation, the uncertainty from the Sherman function is not taken into consideration. The large error bar on one data point in Fig. 3(e) comes from the close values of spin up and spin down counts in the raw data. By definition, identical spin up and down counts give rise to a divergent uncertainty of the spin polarization. Additionally, the positions of $k_{\mathrm{3}}$ and $k_{\mathrm{4}}$ in Fig. 3(f) were actually taken on the opposite sides of the $\bar{K}-\bar{M}-\bar{K^{\prime}}$, closer to the $\bar{\Gamma}$ hexagonal pocket than the cut in Fig. 3(a). We instead showed the cut to be outside of the BZ just for illustration purpose of a complete FS image. The curvature of the momentum cut on the FS in Fig. 3(a) and (f) is omitted to avoid curved and distorted band dispersions in the 3D plot.
ARPES data on InTaS2 and PbTaSe2 were taken at the Stanford Synchrotron Radiation Lightsource, Beamline 5-2 and the Brookhaven National Lab, National Synchrotron Light Source II, Beamline 21-ID, respectively. Both are equipped with a DA30 electron analyzer with vertical slit and have linear horizontal, linear vertical, and circularly polarized photons available. Only the data measured by linear horizontal light were included in the main text and the Supplementary Information. Samples of both kinds were cleaved in-situ under a based pressure better than $3\times 10^{-11}$ Torr and temperatures below 30K. All ARPES measurements in the standard mode maintain an energy and angular resolution superior to 20 meV and 0.1, while spin-ARPES on In1/2TaS2 has the energy and angular resolution better than 50 meV and 1.
@@ 0.7 Ab-initio calculations
The Bloch spectral function and the one-step model ARPES calculations were carried out using the spin-polarized relativistic Korringa-Kohn-Rostoker (SPR-KKR) package [57], under the full-potential fully-relativistic four component Dirac formalism. Exchange-correlation potential within the local spin density approximation by Vosko, Wilk, and Nusair [58] was used. To model the random vacancy on the indium site in InxTaS2 ($x=$ 1/2, 2/3), single-site coherent potential approximation was employed to obtain an auxiliary effective medium that reproduces the concentration averaged scattering properties [59, 57]. The KKR equations were solved with an angular momentum cutoff of $l_{\rm max}$ = 4 to account for the occupied 4f orbitals of Ta and the needs for spectroscopy calculations. The ARPES calculations considered a semi-infinite surface model terminated by the In atoms under the experimental geometry and used the layer-KKR multiple scattering theory, together with coherent potential approximation [60, 61]. Therefore, such theory takes into account all factors such as light polarization, matrix element, final-state, surface, disorder, relativistic, and multiple scattering effects. Experimental crystal structures of In1/2TaS2 and In2/3TaS2 were utilized in the calculations [32, 33].
@@
Data for this study are available in the main text and the Supplementary Information, or can be accessed on Zenodo [70]. Further data that support the findings of this study are available from the corresponding authors upon request.
Code availability
Two ab-initio DFT-based packages were used in this study. The SPR-KKR package is freely available under the specific user license and can be downloaded following registration at https://www.ebert.cup.uni-muenchen.de/index.php/en/software-en. The VASP package can be purchased from https://www.vasp.at/.
@@
M.Y. and E.M. conceived the project. J.H. suggested the material family. Y.Gao synthesized the crystals and performed magnetotransport and magnetization measurements with the help from M.O.,A.M.H., A.Moon, and L.B. under the supervision of E.M.. Y.Z., Y.Guo, Z.Y., and J.S.O. performed the ARPES experiments under the supervision of M.Y. and R.J.B., with the help from A.F., S.K.M., M.H., D.L., A.R., and E.V.. A.P. and Y.Z. performed the KKR calculations, with support from A. Marmodoro, under the supervision of J.M.. X.G. carried out the wave-function based first-principles calculations, with the help from X.J.G., under the supervision of K.T.L.. P.H. theoretically studied topological consequences of general KNLs. Y.Z., Y.Gao, and A.P. contributed equally to this work. All authors contributed to the manuscript preparation.